OpenStudy (anonymous):

Can anyone walk me through the steps of this quadratic equation, tell me what the discriminant is and how to check it for its truth value in the end? a^2 - 6a + 9 = 0

OpenStudy (anonymous):

http://www.bbc.co.uk/bitesize/higher/maths/algebra/quadratic_theory/revision/4/ that is where I learnt it from

OpenStudy (anonymous):

Thank you for the information, looking it over now!

OpenStudy (anonymous):

great. So in your equation a = 1, b = -6 and c = 9 its a shame your variable is a, and the variable in the quadratic equation is also called a. makes it abit confusing

OpenStudy (anonymous):

I know, the quadratic formulas are also difficult for me to write.

OpenStudy (anonymous):

The discriminant is \[ 6^2 - 4(1)9=36-36=0\] So your quadratic is a perfect square \[ a^2 -6 x + 9= (a-3)^2=0\] So the two roots are equal and each one of them is equal to 3

OpenStudy (anonymous):

I should have put \( (-6)^2 \) above.

OpenStudy (anonymous):

Hello eliassaab, thank you for helping me. are you stating that the last part of the equation should be written a&2 (-6) + 9 = (a-3)^2 = 0

OpenStudy (anonymous):

When computing the discriminant \[ b^2 - 4 a c = (-6)^2 - 4(1) 9 = 36-36=0 \]

OpenStudy (anonymous):

It does not matter since \[(-6)^2 = 6^2 =36\]

OpenStudy (anonymous):

I was just confirming what you wrote. Thank you!

OpenStudy (anonymous):

YW

OpenStudy (jhannybean):

Yep, when the discriminant \(\large D = 0\) , you result in one real zero (answer) with a multiplicity of 2.

OpenStudy (anonymous):

Thanks Jhannybean. I am trying to write the answer step by step.

OpenStudy (jhannybean):

Good luck! :) And no problem.

OpenStudy (anonymous):

Thank you!